# MFCLIP: Multi-modal Fine-grained CLIP for Generalizable Diffusion Face Forgery Detection

### v1

@@ IV-B Comparison with the State of the Art
Cross-generator evaluation. To conduct an in-depth study of the proposed network for generalizable DFFD, we performed the cross-diffusion evaluation on GenFace. Since DFFD is a new and challenging task, to the best of our knowledge, there have been no comprehensive experiments to evaluate the performance of FFD models on cross-diffusion generators. Therefore, our work is the first to assess the generalization of models to unseen diffusion-generated facial images, systematically and comprehensively. Specifically, we trained models using the images generated by one diffusion-based generator and tested them on different ones. As Table III shows, detectors tend to achieve better performance on images produced by the generator with high similarity to the generator used for training. Specifically, since DDPM and DiffFace resemble each other, the detector trained using DDPM shows excellent results on images synthesized by DiffFace, and vice versa. For instance, the AUC of Xception, CViT and DIRE is 96.80%, 96.21%, and 94.33% on DiffFace after training using DDPM, respectively. By contrast, they only acquire 63.57% AUC, 74.51% AUC, and 62.50% AUC on LatDiff, individually.
@@
In addition, as Fig. 6 displays, networks generally perform worse on the images generated by LatDiff than that synthesized by other generators such as DDPM, CollDiff, DiffFace, and Diffae. We believed that facial images produced by LatDiff are realistic, such that the detector struggles to distinguish their authenticity. Therefore, those generated by LatDiff pose a huge threat to the detector. We further performed the cross-generator protocol on AM and FS, respectively. As Table IV and Table II display, most models acquire poor performance (about 60% AUC) on the cross-generator evaluation. The AUC of our network is around 40.60%, 27.26%, and 26.36% higher than that of CLIP, DIRE, and FreqNet, respectively, on IAFaces after training using Diffae, which is attributed to the powerful modelling capability of our MFCLIP method.
@@
Cross-dataset evaluation. To investigate the generalization of MFCLIP, we conducted cross-dataset evaluations. We trained detectors using FF++ and tested them on FF++, CelebDF, DFDC, and DF-1.0. The first two level text prompts in FF++ are only introduced to train MFCLIP, due to the limitation of labels offered by FF++. As Table V shows, the AUC of our MFCLIP method is about 1.34% higher than that of VLFFD, showing the superior generalization ability of MFCLIP.
@@
Robustness to common image corruptions. We assessed the robustness of detectors against different unseen image distortions. We trained models on GenFace and tested their performance on distorted images from [32]. Seven types of perturbations are involved, each with five intensity levels. As shown in Fig. 7, we tested the models on various image distortions, such as saturation changes, contrast adjustments, block distortions, white Gaussian noise, blurring, pixelation, and video compression. We averaged the AUC scores of the detector on seven types of corrupted images. An intensity of 0 indicates no degradation. When adopting perturbations of different severities, the changes in AUC for all detectors are presented in Fig. 7. The results show that our model consistently surpasses all other methods across seven types of image degradation.
@@
Figure 8: (a) and (b) show the ablation results of CLIP and MFCLIP with hierarchical fine-grained text prompts, respectively.
@@
TABLE VI: MFCLIP ablation. We tested models on DDPM, LatDiff, DiffFace, and Diffae, after training using CollDiff.
@@ V Ablation Study
We conducted an ablation study to evaluate the contribution of each component in MFCLIP. In this study, we used images generated by five diffusion-based models including CollDiff, DDPM, LatDiff, DiffFace and Diffae, to investigate the effectiveness of MFCLIP. We considered five schemes: 1) the impacts of components, 2) the influence of sample pair attention, 3) the effects of fine-grained text prompts, 4) the influences of loss functions and 5) the effect of the patch size in PS. In the subsequent subsection, we discussed the five aspects, respectively.
@@ V-A Impacts of components
To examine the contribution of each component to learning ability, we observed the performance of models on DDPM, LatDiff, DiffFace, and Diffae after training using CollDiff. Table VI shows the ablation results of the model. NE improves the performance by 47.89% AUC on LatDiff, confirming that noises extracted from the richest patch offer valuable information to benefit the DFFD. The gains from introducing the FLE module (+45.55%) are obvious, demonstrating the significance of fine-grained text embeddings. Predictor increases performance by 3.23% ACC on DiffFace, showing that DFFD could benefit from aligning visual and linguistic representations in feature space. The introduction of SPA further enhances the performance (+3.79%), which verifies that cross-modality sample pairs are adaptively emphasized and suppressed, guiding the model to achieve better feature alignment.
TABLE VII: Ablation results of various patch sizes. We tested models on DDPM, LatDiff, DiffFace, and Diffae, after training on CollDiff.
@@ V-B Effect of the patch size in PS
We investigated the impact of different patch sizes in PS. The proposed MFCLIP model was trained using CollDiff and tested on DDPM, LatDiff, DiffFace, and Diffae. We reported the performance of MFCLIP from patch size 16 to 224. Table VII shows that the performance typically increases with the growth of patch size. The AUC reaches the peak as the patch size is 112, and begins to decline afterward. We argued that the model tends to explore more forgery areas, when the larger patch size is leveraged. However, oversized patch scales like 224 may introduce noises, enabling the model to acquire poor generalization performance.
TABLE VIII: The effect of the SPA module. We tested models on DDPM, LatDiff, and Diffae, after training using CollDiff.
@@ V-C Influence of sample pair attention
To evaluate the effectiveness and efficiency of SPA, we conducted the cross-generator evaluation to show that our plug-and-play SPA module can generalize to various vison-language-based models for performance improvement. We trained models using CollDiff and tested them on DDPM, LatDiff, and Diffae. As Table VIII shows, due to the addition of MAS, the AUC of both CLIP and MFCLIP is improved by 2.72% and 3.78%, respectively, on the challenging LatDiff. Furthermore, SPA rarely introduces auxiliary weights and computational costs when plugged and played into other models.
@@ V-D Effects of fine-grained text prompts
To investigate the impact of the hierarchical fine-grained text prompts, we evaluated the performance of CLIP and our MFCLIP method on DDPM, LatDiff, and Diffae after training on CollDiff, by gradually introducing the hierarchical texts. In Fig. 8, we noticed that the performance of our model commonly improves with the increase of hierarchical text prompts. Specifically, the AUC of CLIP and our model is improved on DDPM, LatDiff, and Diffae, respectively, as fine-grained texts from level 1 to level 4 are introduced, showing that hierarchical text prompts guide the model to capture more general and discriminative information, to facilitate the advancement of DFFD.
TABLE IX: Effects of losses. We tested models on DDPM, LatDiff, DiffFace, and Diffae, after training on CollDiff.
@@
V-E Influences of loss functions
@@
To verify the contribution of loss functions, we performed ablations on various losses. The ablation result of losses is displayed in Table IX. As MFCLIP is guided with merely cross-entropy loss, the AUC is 95.48% on LatDiff, but an about 1.8% increase of AUC could be reached via adding the KL loss. We believed that it could enhance the visual forgery representations via language guidance. Meanwhile, due to the introduction of the CMC loss, MFCLIP is grown by 2.6% AUC on LatDiff, showing the effectiveness of adaptive cross-modal pairs alignment. The integration of the three losses performs the best among these losses, which suggests that the proposed loss could acquire promising results.

### v2

@@ IV-B1 Cross-generator evaluation
To conduct an in-depth study of the proposed network for generalizable DFFD, we perform the cross-diffusion evaluation on GenFace. Since DFFD is a new and challenging task, to the best of our knowledge, there have been no comprehensive experiments to evaluate the performance of FFD models on cross-diffusion generators. Therefore, our work is the first to assess the generalization of models to unseen diffusion-generated facial images, systematically and comprehensively. Specifically, we train models using the images generated by one diffusion-based generator and test them on different ones. As Table III shows, detectors tend to achieve better performance on images produced by the generator with high similarity to the generator used for training. Specifically, since DDPM and DiffFace resemble each other, the detector trained using DDPM shows excellent results on images synthesized by DiffFace, and vice versa. For instance, the AUC of Xception, CViT and DIRE is 96.80%, 96.21%, and 94.33% on DiffFace after training using DDPM, respectively. By contrast, they only acquire 63.57% AUC, 74.51% AUC, and 62.50% AUC on LatDiff, individually.
@@
In addition, as Fig. 7 (c) displays, networks generally perform worse on the images generated by LatDiff than on that synthesized by other generators such as DDPM, CollDiff, DiffFace, and Diffae. We believe that facial images produced by LatDiff are realistic, such that the detector struggles to distinguish their authenticity. Therefore, those generated by LatDiff pose a huge threat to the detector. We further perform the cross-generator protocol on AM and FS, respectively. As Table IV and Table II display, most models acquire poor performance (about 60% AUC) on the cross-generator evaluation. The AUC of our network is around 40.60%, 27.26%, and 26.36% higher than that of CLIP, DIRE, and FreqNet, respectively, on IAFaces after training using Diffae. Besides, the AUC of our MFCLIP model is about 24% higher than that of CAEL on DDPM after training using CollDiff. We believe that our MFCLIP model can extract more fine-grained language representations and conduct vision-language contrastive learning, adaptively, thus obtaining excellent performance.
IV-B2 Cross-forgery evaluation
To investigate the generalization of various detectors, we perform cross-forgery tests. We train models using images of one manipulation, and test them on those of remaining manipulations. As Table I shows, the performance of our methods outperforms most detectors, demonstrating the superior generalization of MFCLIP. Specifically, for the vision-language-based models, the AUC of our model is nearly 26.80%, 18.78%, and 14.66% higher than that of CLIP, VLFFD, and DD-VQA, respectively, on EFS after training using AM, which is attributed that our MFCLIP model not only captures global visual forgery patterns across image noise modalities but also achieves better vision-language alignment.
IV-B3 Cross-dataset evaluation
@@
To investigate the generalization of MFCLIP, we conduct cross-dataset evaluations. We train detectors using FF++ and test them on FF++, CelebDF, DFDC, and DF-1.0. The first two level text prompts in FF++ are only introduced to train MFCLIP, due to the limitation of labels offered by FF++. As Table V shows, the AUC of our MFCLIP method is about 1.34% higher than that of VLFFD. We argue that our MFCLIP model can capture discriminative global noise features, and achieve more flexible vision-language alignment, compared to the VLFFD model.
IV-B4 Generalization to unseen generative models
To further study the generalization of different detectors to unseen generative models, we train networks using images generated by various generators from GenFace, and test them on those synthesized by autoregressive, VAE-based, and flow-based approaches. In particular, we leverage autoregressive models (e.g. HART [58] and VAR [6]), VAE-based generators (e.g. vanilla VAE [1]), and flow-based generative networks (e.g. FLUX [59] and FlowEdit [8]), to generate 2k face manipulated images, respectively. We also choose 2k fake face images created by the VAE-based DF-VAE [39] model from the DF-1.0 dataset. We evaluate the performance of various detectors on these face forgery images, accordingly. In Table IX, methods tend to achieve lower ACC scores on face images created by VAR than on those generated by HART. This indicates that VAR which shows SOTA performance creates high-quality face images that are more complex and challenging. To further improve generalization to different forgery types, we introduce various generator distributions under level 4 in GenFace to train the detector, to capture more general forgery features. Specifically, we arbitrarily adopt 45k real face images and 45k fake ones from GenFace, with the fake images evenly distributed across ten different GAN or diffusion generators under level 4 (4.5k images per generator). To ensure a fair comparison, all detectors are trained using the same generator distributions from GenFace. In Table IX, the AUC of MFCLIP is around 88% on flow-based models. Compared to the state-of-the-art DIRE, MFCLIP achieves superior performance on images created by flow-based generators.
Despite the better generalization to different forgery types than other detectors, we need to improve transferability across diverse generators such as autoregressive, VAE-based, and flow-based methods. In the future, we aim to explore discriminative prior characteristics of various forgery types, such as noise patterns in flow-based methods or reconstruction errors in autoregressive models, and incorporate them into the detection model. Besides, multi-modal large language models could be integrated with our detector to capture more complex, nuanced, and general patterns in forgery face data.
IV-B5 Robustness to common image corruptions
@@
We assess the robustness of detectors against different unseen image distortions. We train models on GenFace and test their performance on distorted images from [39]. Ten types of perturbations are involved, each with five intensity levels. As shown in Fig. 6, we test the models on various image distortions, such as saturation changes, contrast adjustments, block distortions, white Gaussian noise, blurring, pixelation, video compression, upscaling, downscaling, and cropping. An intensity of 0 indicates no degradation.
Specifically, the specific settings for the severity of the first seven deformations are discussed in detail in [40]. Upscaling aims to enlarge the image pixel dimensions using scaling factors [1.1, 1.3, 1.5, 2.0, 3.0] with five intensities (the larger the factor value, the greater the upscaling rate). Downscaling intends to reduce the image pixel dimensions by applying scaling parameters such as [0.91, 0.77, 0.67, 0.50, 0.33], which correspond to five intensities of downscaling (the smaller the parameter value, the greater the downscaling degree). Cropping removes a portion of the image based on crop percentages [0.8, 0.7, 0.6, 0.5, 0.4], with each percentage representing a different level of image retention (the smaller the percentage, the greater the cropping severity). When adopting perturbations of different severities, the changes in AUC for all detectors are presented in Fig. 6. We observe that the AUC of models generally decreases with the increase of downscaling intensities. By contrast, the AUC of our MFCLIP method shows an increased trend, which demonstrates the strong robustness of MFCLIP. Besides, the performance of detectors reaches the minimum when the severity of cropping equals one, and begins to rise thereafter. This trend indicates that detectors could generalize well to smaller facial areas when trained using whole face images. The results show that our model surpasses most methods across various types of image degradation.
@@
Figure 7: (a) and (b) display the ablation results of CLIP and MFCLIP with hierarchical fine-grained text prompts, respectively. (c) shows the comparison of the performance of detectors on various generators. We average the AUC scores of the model listed in Table III on the images from the same generator.
@@
TABLE VI: MFCLIP ablation. We test models on DDPM, LatDiff, DiffFace, and Diffae, after training using CollDiff in GenFace.
@@ V Ablation Study
We conduct an ablation study to evaluate the contribution of each component in MFCLIP. In this study, we use images generated by five diffusion-based models including CollDiff, DDPM, LatDiff, DiffFace and Diffae, to investigate the effectiveness of MFCLIP. We consider six schemes: 1) the impacts of components, 2) the influence of sample pair attention, 3) the effects of fine-grained text prompts, 4) the influences of loss functions, 5) the effect of the patch size in PS, 6) Impacts of pre-trained weights of CLIP, and 7) Impacts of the FID distribution in FTG. In the subsequent subsection, we discuss the seven aspects, respectively.
TABLE VII: The FID scores of various forgery categories under different levels in FTG. We compute the FID scores on the GenFace dataset. The lower FID scores indicate better image quality.
L1 L2 L3
Fake EFS FS AM Average EFS GAN EFS Diffusion FS GAN FS Diffusion AM GAN AM Diffusion Average
FID 13.097 13.353 17.145 31.257 20.585 40.041 12.368 21.556 13.908 44.349 42.253 29.079
L4
StyleGAN2 StyleGAN3 CollDiff DDPM LatDiff FSLSD FaceSwapper DiffFace Diffae LatTrans Average
FID 39.102 42.654 20.855 35.738 11.274 6.517 50.738 13.908 42.253 44.349 30.739
@@ V-A Impacts of components
To examine the contribution of each component to learning ability, we observe the performance of models on DDPM, LatDiff, DiffFace, and Diffae after training using CollDiff. Table VI shows the ablation results of the model. NE improves the performance by 47.89% AUC on LatDiff, confirming that noises extracted from the richest patch offer valuable information to benefit the DFFD. The gains from introducing the FLE module (+45.55%) are obvious, demonstrating the significance of fine-grained text embeddings. Predictor increases performance by 3.23% ACC on DiffFace, showing that DFFD could benefit from aligning visual and linguistic representations in feature space. The introduction of SPA further enhances the performance (+3.79%), which verifies that cross-modality sample pairs are adaptively emphasized and suppressed, guiding the model to achieve better feature alignment.
TABLE VIII: Ablation results of various patch sizes. We test models on DDPM, LatDiff, DiffFace, and Diffae, after training on CollDiff in GenFace.
@@
TABLE IX: Generalization to various generative models. We test detectors on autoregressive, VAE-based, and flow-based generative methods after training using GenFace.
Method Autoregressive VAE-based Flow-based
HART VAR
ACC AUC ACC AUC ACC AUC ACC AUC
CViT [14] 55.50 87.87 54.08 83.88 50.13 65.79 56.05 84.31
FreqNet [17] 56.10 84.52 52.75 77.41 52.63 72.82 62.72 87.75
DIRE [13] 61.87 87.69 58.26 85.94 57.64 76.80 62.49 85.76
CAEL [40] 54.55 83.77 56.99 81.14 53.62 76.67 61.91 86.04
MFCLIP 62.74 88.24 60.32 86.35 60.94 77.31 64.90 88.19
@@ V-B Effect of the patch size in PS
We investigate the impact of different patch sizes in PS. The proposed MFCLIP model is trained using CollDiff and tested on DDPM, LatDiff, DiffFace, and Diffae. We report the performance of MFCLIP from patch size 16 to 224. Table VIII shows that the performance typically increases with the growth of patch size. The AUC reaches the peak as the patch size is 112, and begins to decline afterward. We argue that the model tends to explore more forgery areas, when the larger patch size is leveraged. However, oversized patch scales like 224 may introduce noises, enabling the model to acquire poor generalization performance.
TABLE X: The effect of the SPA module. We test models on DDPM, LatDiff, and Diffae, after training using CollDiff in GenFace.
@@ V-C Influence of sample pair attention
To evaluate the effectiveness and efficiency of SPA, we conduct the cross-generator evaluation to show that our plug-and-play SPA module can generalize to various vison-language-based models for performance improvement. We train models using CollDiff and test them on DDPM, LatDiff, and Diffae. As Table X shows, due to the addition of SPA, the AUC of both CLIP and MFCLIP is improved by 2.72% and 3.78%, respectively, on the challenging LatDiff, with almost no change in the number of parameters and FLOPs. SPA rarely introduces auxiliary weights and computational costs when plugged and played into other models like CLIP. This shows that performance improvements could be achieved with high computational efficiency as the addition of SPA.
TABLE XI: The impact of the fine-grained text prompts. * denotes wider text prompts in the hierarchical structure. L5 means that we introduce the learnable text prompts to build the deeper hierarchical structure. We test models on DDPM, LatDiff, and Diffae, after training using CollDiff in GenFace.
Method DDPM LatDiff Diffae Params (M) FLOPs (G)
ACC AUC ACC AUC ACC AUC
Ours w/L1 99.01 99.04 95.82 96.03 90.45 97.05 93.83 218.60
Ours w/L1+L2 99.25 99.38 96.38 96.57 90.72 97.34 93.83 265.11
Ours w/L1+L2+L3 99.66 99.79 97.03 97.36 91.08 97.63 93.83 311.62
Ours w/L1+L2+L3+L4 100 100 99.19 99.97 93.94 99.99 93.83 358.12
Ours w/L1*+L2+L3+L4 99.98 100 96.48 99.79 95.82 99.57 93.83 358.12
Ours w/L1+L2+L3+L4+L5 100 100 98.92 99.66 98.44 99.40 93.83 404.63
@@ V-D Effects of fine-grained text prompts
To investigate the impact of the hierarchical fine-grained text prompts, we evaluate the performance of CLIP and our MFCLIP method on DDPM, LatDiff, and Diffae after training on CollDiff, by gradually introducing the hierarchical texts. In Fig. 7, we notice that the performance of our model commonly improves with the increase of hierarchical text prompts. Specifically, the AUC of CLIP and our model is improved on DDPM, LatDiff, and Diffae, respectively, as fine-grained texts from levels 1 to 4 are introduced, showing that hierarchical text prompts guide the model to capture more general and discriminative information, to facilitate the advancement of DFFD.
Additionally, Table XI shows the performance of MFCLIP with wider or deeper text prompts in the hierarchical structure. In detail, at level 1, we generate detailed text prompts such as “a photo of a real face with natural skin texture, facial symmetry with natural variation, life-like expressions, and dynamic facial contours” or “a photo of a fake face with unrealistic skin texture, perfect symmetry without variation, unnatural facial proportions, lack of life-like expressions, inconsistent lighting and shadows, and a flat, two-dimensional appearance”, to create a wider hierarchical structure. We also formulate learnable text prompts at level 5, to build a deeper one. We observe that using deeper or wider text prompts in the hierarchical structure struggles to improve the performance of MFCLIP, as they lead to information redundancy, weakening the representation capability. Besides, MFCLIP achieves excellent results while maintaining the trade-off between computational efficiency and performance, when four levels of text prompts are involved. Specifically, an approximately 2% increase of AUC could be achieved by adding a level of text prompts from level 3 to 4, with no change in the number of parameters but an increase of 47.5G FLOPs. The AUC shows a decreased trend when the level of text prompts equals five.
TABLE XII: Effects of losses. We test models on DDPM, LatDiff, DiffFace, and Diffae, after training on CollDiff in GenFace.
@@
TABLE XIII: Effects of various generator distributions. We test models on DDPM, DiffFace, and Diffae, after training on LatDiff or both LatDiff and StyleGAN3. SG3 is StyleGAN3.
Training Set Model Testing Set
DDPM CollDiff DiffFace Diffae
ACC AUC ACC AUC ACC AUC ACC AUC
LatDiff Xception [11] 50.13 76.45 50.00 37.05 51.51 96.21 50.28 92.62
LatDiff+SG3 Xception [11] 51.51 94.23 50.14 36.27 53.64 98.68 64.77 98.34
LatDiff CLIP [26] 62.33 74.53 52.99 53.23 55.71 61.05 56.64 64.38
LatDiff+SG3 CLIP [26] 65.02 83.65 50.58 48.29 55.96 68.04 59.45 68.76
LatDiff MFCLIP 99.99 99.99 65.08 77.07 99.98 99.98 97.92 99.99
LatDiff+SG3 MFCLIP 100 100 65.10 77.84 99.99 99.99 98.32 100
V-E Impacts of the FID distribution in FTG.
To study the correlation between the FID and the criteria for level classifications in FTG, we calculate the FID score of various generators or manipulations at different levels in FTG. To guarantee a fair and balanced assessment, we keep an equivalent distribution of pristine and forgery face images, i.e., a 1:1 ratio. In detail, for level 1, we randomly select 10k real face images and 10k fake ones, with the fake ones evenly distributed across ten deepfake generators (1k images per generator). Similarly, at levels 2 and 3, we choose an equal number of real and fake images to calculate the FID scores for each forgery, respectively. At level 4, we also adopt 10k real images and 10k fake ones to compute the FID score of each generator.
The results across various levels in FTG are displayed in Table VII. We observe that the difference in FID scores across categories under each level becomes increasingly pronounced, as the hierarchy goes deeper. This phenomenon suggests that classification may become easier, as the number of categories increases from 2 (L1) to 4 (L2), 7 (L3), and 11 (L4) [40]. Moreover, the FID scores exhibit a corresponding rise from 13.097 (L1) to 20.585 (L2), 29.079 (L3), and 30.739 (L4), with the hierarchy level increasing, which facilitates the deepfake detection since there is a growing distinction between authentic and forged images. It is observed that the FID score of the diffusion mode is lower than that of GAN under different forgeries in L3, indicating that diffusion face images achieve superior quality. Therefore, we argue that the diffusion mode is inclined to push detectors to study more discriminative and general forgery artifacts, thus improving generalization capabilities.
We notice that when the distribution is the most different from the others in GAN data, the FID score is improved. To study the impact of different distributions on model performance, face images created by LatDiff and StyleGAN3 with the largest distribution (i.e. FID) difference under the EFS forgery are selected to train detectors. In Table XIII, the AUC of the detector trained using both LatDiff and StyleGAN3 outperforms that of the model trained with LatDiff. This suggests that the performance of detectors tends to increase, when facial images with significant distributional differences are introduced. We argue that the new distribution may increase the diversity of data, allowing the model to learn more robust features, and thereby improving generalization capabilities.
V-F Influences of loss functions
@@
To verify the contribution of loss functions, we perform ablations on various losses. The ablation result of losses is displayed in Table XII. As MFCLIP is guided with merely cross-entropy loss, the AUC is 95.48% on LatDiff, but an about 1.8% increase of AUC could be reached via adding the KL loss. We believe that it could enhance the visual forgery representations via language guidance. Meanwhile, due to the introduction of the CMC loss, MFCLIP is grown by 2.6% AUC on LatDiff, showing the effectiveness of adaptive cross-modal pairs alignment. The integration of the three losses performs the best among these losses, which suggests that the proposed loss could acquire promising results.
TABLE XIV: Effects of pre-trained weights of CLIP. † denotes the image encoder followed by the image adapter or image LoRA. * denotes the text encoder followed by a text adapter or text LoRA. ‡ denotes both image and text encoders, each followed by an adapter or LoRA. We test models on DDPM, LatDiff, DiffFace, and Diffae, after training on CollDiff in GenFace. FFT means full fine-tuning.
Model DDPM LatDiff DiffFace Diffae
ACC AUC ACC AUC ACC AUC ACC AUC
CLIP 51.58 53.92 50.21 46.17 50.26 48.64 49.82 48.36
CLIP-FFT 62.38 84.01 50.02 56.39 51.91 70.28 50.51 60.62
CLIP-Adapter† 70.66 78.06 48.32 45.06 59.32 70.74 55.40 62.51
CLIP-Adapter* 63.56 70.38 48.06 44.30 65.16 73.04 58.19 57.49
CLIP-Adapter‡ 65.10 80.27 48.79 43.45 63.48 72.32 64.97 67.76
CLIP-LoRA† 80.52 84.66 50.78 49.07 68.64 75.51 70.98 75.11
CLIP-LoRA* 81.87 84.93 46.97 44.20 68.02 75.54 67.54 74.37
CLIP-LoRA‡ 80.36 84.63 48.67 45.02 68.49 75.08 68.96 73.01
MFCLIP 100 100 99.19 99.97 99.63 99.96 93.94 99.99
V-G Impacts of pre-trained weights of CLIP
To study the effect of pre-trained weights of CLIP, we fine-tune pre-trained CLIP models using face images generated by CollDiff, and test them on DDPM, LatDiff, DiffFace, and Diffae. Specifically, we adjust all parameters of the pre-trained CLIP model via full fine-tuning. To conduct parameter-efficient fine-tuning, we incorporate residual-style adapters or rank decomposition matrices at the end of CLIP’s image and text encoder, respectively, following CLIP-Adapter [60] and LoRA [61] (rank=8, alpha=32). We freeze pre-trained CLIP’s image and text encoders, and fine-tune using either the image adapter, image LoRA, text adapter, or text LoRA, accordingly. We also fine-tune CLIP using both the text and image adapters or both the text and image LoRA. In Table XIV, it is noticed that CLIP fine-tuned with LoRA achieves the best performance among various fine-tuned CLIPs. Besides, we can observe that the performance of CLIP with the updated image adapter outperforms that of CLIP with the modified text adapter. We argue that image feature adaptation is more important than text feature adjustment for DFFD, since the domain gap of visual embeddings between the pre-trained dataset and face forgery dataset is bigger than that of text representations.
Besides, we train the vanilla CLIP from scratch without loading pre-trained weights. Table XIV shows that the performance of fine-tuned CLIPs is higher than that of the CLIP trained from scratch. We believe that fine-tuning enables CLIP to effectively utilize the feature representations learned during pre-training on large-scale datasets. This allows for improved generalization, particularly for tasks involving subtle distinctions, such as FFD. The performance of our MFCLIP method outperforms that of fine-tuned CLIP models, demonstrating that both global image-noise fusion and flexible vision-language matching are helpful for DFFD.